Circular No. 7045

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1998 U5 (LINEAR) Preliminary parabolic orbital elements from 44 observations (all of

which appear on MPEC 1998-V03), spanning Oct. 30–Nov. 2:

T = 1998 Dec. 21.774 TT  ω = 51.^{o}448
Ω = 66.661 }2000.0
q = 1.23192 AU  i = 131.999
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 30 6^{h}54.^{m}92 +28^{o}28.^{′}5 0.753 1.471 114.^{o}2 38.^{o}0 12.1
Nov. 1 6 47.02 +31 16.8 0.697 1.455 117.9 37.1 11.8
3 6 36.84 +34 31.2 0.643 1.440 121.9 35.8 11.6
5 6 23.50 +38 14.8 0.593 1.425 126.1 34.2 11.4
7 6 05.66 +42 28.2 0.548 1.410 130.4 32.4 11.2
9 5 41.27 +47 06.9 0.509 1.396 134.4 30.5 11.0
11 5 07.37 +51 55.8 0.478 1.382 137.7 28.8 10.8
13 4 20.30 +56 21.9 0.456 1.368 139.4 28.1 10.7
15 3 18.12 +59 31.9 0.444 1.355 138.8 28.7 10.6
17 2 05.92 +60 29.7 0.443 1.343 135.9 30.8 10.5
19 0 56.74 +59 00.1 0.454 1.331 131.0 34.1 10.5
21 0 00.92 +55 43.9 0.475 1.320 125.2 37.7 10.6
23 23 19.85 +51 38.9 0.505 1.309 119.1 41.2 10.7

Total visual magnitude and coma-diameter estimates: Nov. 1.18 UT, 12.3, 1′.3 (J. D. Shanklin, Cambridge, England, 0.30-m refractor); Nov. 1.51, 12.2, 2′ (A. Hale, Cloudcroft, NM, 0.41-m reflector; poor conditions).

SUPERNOVA 1998en IN UGC 3645 W. D. Li, M. Modjaz, E. Halderson, T. Shefler, J. Y. King, M. Pa-

penkova, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report the discovery of an apparent supernova in the course of the Lick Observatory Supernova Search with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1998en was discovered on unfiltered frames taken on Oct. 30.5 (mag ∼ 18.4) and 31.5 UT (mag ∼ 18.2). The new object is located at α = 7^{h}04^{m}00^{s}.05, δ = +50^{o}40′30′′.0 (equinox 2000.0), which is ∼ 10′′.4 west and 18′′.3 south of the nucleus of UGC 3645. KAIT images of the same field on Oct. 26.5 (limiting mag ∼ 18.5) and Oct. 21.5 (limiting mag ∼ 19.0) showed nothing at the position of the supernova.

1998 November 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green